OpenStudy (anonymous):

Which graph represents the function of f(x) = the quantity of 4 x squared minus 4 x minus 8, all over 2 x plus 2? A.graph of 2 x minus 4, with discontinuity at negative 1, negative 6 B.graph of 2 x minus 4, with discontinuity at 1, negative 2 C.graph of 2 x plus 2, with discontinuity at negative 1, 0 D.graph of 2 x plus 2, with discontinuity at 1, 4

OpenStudy (anonymous):

@satellite73 @princeharryyy @FEARLESS_JOCEY @welshfella @Etrainx

OpenStudy (welshfella):

4x^2 - 4x - 8 ------------ 2x + 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait what

OpenStudy (welshfella):

this simplifies to a linear function

OpenStudy (welshfella):

so its graph will be a straight line

OpenStudy (welshfella):

it simplifies to 4 ( x + 1)(x - 2) ------------- 2(x + 1) = 4(x - 2) / 2 = 2(x - 2)

OpenStudy (welshfella):

= 2x - 4 where will the discontinuities be?

OpenStudy (welshfella):

- sorry theres just one discontinuity

OpenStudy (welshfella):

although it simplifies to a linear function and the graph will be a line except, because it is a fraction there will be a point where there is no value for f(x) and it can be represented on the graph by a small open circle.

OpenStudy (welshfella):

- this is called a 'hole' in the graph This one is confusing me a little because the way to find holes is to set the factor common to top and bottom of the fraction to zero and solve for x in this case the factor is x + 1 x + 1 = 0 so x = -1 that is a hole at x = -1

OpenStudy (welshfella):

and corresponding value of f(x) = -6 oh OK - i got confused That is the answer B

OpenStudy (welshfella):

drat I meant A